OpenStudy (anonymous):

If both the base and height of a triangle are made one hundred times longer, how many times greater is the new area?

OpenStudy (anonymous):

Area of triangle is given by: \[Area = \frac{1}{2} \times Base \times Height\] Now Base and Height are made 100 times longer that is : Base = 100 Base and Height = 100 Height then: \[New \;\; Area = \frac{1}{2} \times (100 Base) \times (100 Height)\] \[New \;\; Area = 10000 \times \color{blue}{\frac{1}{2} \times Base \times Height}\] \[\color{green}{New \;\; Area = 10000 \times (Original \;\;Area)}\]

OpenStudy (anonymous):

so it is ten thousand times more? @waterineyes

OpenStudy (anonymous):

Yes it is..

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Welcome dear..